# Predictive Non-linear Dynamics via Neural Networks and Recurrence Plots

### v1

@@ Abstract
Predicting and characterizing diverse non-linear behaviors in dynamical systems is a complex challenge, especially due to the inherently presence of chaotic dynamics. Current forecasting methods are reliant on system-specific knowledge or heavily parameterized models, which can be associated with a variety of drawbacks including critical model assumptions, uncertainties in their estimated input hyperparameters, and also being computationally intensive. Moreover, even when combined with recurrence analyses, these approaches are typically constrained to chaos identification, rather than parameter inference. In this work, we address these challenges by proposing a methodology that uses recurrence plots to train convolutional neural networks with the task of estimating the defining control parameters of two distinct non-linear systems: (i) the Logistic map and (ii) the Standard map. By focusing on the neural networks’ ability to recognize patterns within recurrence plots, we demonstrate accurate parameter recovery, achieving fairly confident levels of prediction for both systems. This method not only provides a robust approach to predicting diverse non-linear dynamics but also opens up new possibilities for the automated characterization of similar non-linear dynamical systems.
@@ I Introduction
Making accurate predictions in the field of nonlinear dynamical systems can be a challenging task because the evolution of nonlinear relations gives rise to different types of dynamical behavior [1, 2]. Leaving aside induced or inherently stochastic effects, these behaviors can include periodic, quasi-periodic and chaotic dynamics [3]. Periodic motion follows well-defined and regular cycles, whereas quasi-periodic dynamics lack a fixed period but still exhibit ordered motion [4, 5]. Chaotic dynamics, characterized by extreme sensitivity to initial conditions (ICs), evolves unpredictably despite being governed by deterministic rules [6, 7]. Identifying, understanding and especially characterizing this behavior is essential to gain deeper insights into the fundamental nature and properties of dynamical systems of interest.
@@
Parallel to that, developments in the field of neural networks have flourished in the last decade [8], with the introduction of several robust and versatile architectures that allow accurate identification of images in classification problems. Recently, the relevance of neural networks has been highlighted by the 2024 Nobel Prize in Physics, awarded to Geoffrey E. Hinton and John J. Hopfield. Convolutional Neural Networks (CNN) [9, 10] are a subset of feed-forward neural networks that are widely used in image, text and audio recognition tasks, providing a strong foundation for natural language processing techniques that are now state of the art in artificial intelligence, while remaining relevant due to their effectiveness in many areas. In nonlinear dynamical systems, CNNs have also been shown to identify chaotic dynamics when combined with recurrence analysis, achieving over 92% accuracy [11, 12].
Alongside these neural network advancements, Recurrence Plots (RPs) have become a well-established approach in nonlinear data analysis for visualizing the recurrence of states in dynamical systems [13, 14, 15]. Different types of dynamical behavior, such as regular, chaotic or even stochastic motion, produce different visual recurrence patterns. [16, 17] and RPs are essentially images composed of white and black pixels representing non-recurrent and recurrent state, respectively, making them particularly suitable inputs for CNNs. These neural networks can then be trained to identify patterns that correspond to the underlying dynamics, offering an automated process for the characterization of different dynamical behaviors. While RPs have been widely used to detect and study chaotic dynamics [18, 19, 20, 21], their potential for parameter estimation remains relatively unexplored. By encoding a system’s dynamics into a structured visual format, RPs provide a direct method for analyzing behavior and, when combined with CNNs, open up possibilities for inferring system control parameters from recurrence patterns.
In this paper, we propose a methodology that uses CNNs to both classify dynamic behavior and recover control parameters in both one- and two-dimensional nonlinear systems, specifically the logistic and standard maps. By training a CNN on the RPs of these systems, the neural network can predict key control parameters with a fairly high degree of accuracy. Furthermore, the same framework is able to effectively classify different dynamic regimes within the logistic map. Our results demonstrate that CNNs combined with RPs can serve as a powerful numerical approach for characterizing system behavior and identifying the specific parameters that govern it, providing a novel way to apply deep learning to the study of nonlinear dynamics.
The rest of this paper is organized as follows. Section II discusses both the methodology and the models in detail, including the construction of the RPs, the logistic and standard maps and their range of dynamic behaviors, and the architecture of the CNN. Section III presents the results, analyzing the performance of the CNN in both classification and parameter prediction tasks, and discussing the broader implications of these results for the study of nonlinear systems. Finally, section IV concludes the paper with a summary of the results and suggestions for future work, including potential improvements and applications to real-world systems.
@@ II.1 Recurrence Plots
The essence of the proposed methodology relies on the generation of many distinct recurrence plots (RPs), which serve as the fundamental visual representation of the different non-linear dynamical behaviors of interest. It is common to define recurrence if at time $t_{j}$, a given trajectory $\bm{x}(t_{j})\approx\bm{x}(t_{i})$ approximately returns to a previous state $\bm{\varphi}(t_{i})$, with $t_{i}<t_{j}$, meaning it returns to a neighborhood of a past state. Since it is a finite-time analysis, it is worth remarking that $0<t_{i}<t_{j}\leq T$, where $T$ is the trajectory’s maximum iteration time.
Depending on the model, its control parameters and initial conditions, different dynamical behaviors may appear and all respective RPs should be constructed considering these differences. Consequently, all RPs were built with a fixed recurrence rate $(RR)$ i.e., the percentage of recurrent points [14], allowing for an unbiased and appropriate comparison among chaotic, periodic, and quasi-periodic dynamics.
@@
Once the best suited $\varepsilon$ is determined, the binary Recurrence Matrix (RM) $R\in\{0,1\}^{T\times T}$, composed by the elements $R_{i,j}$ can be computed in terms of the distance matrix $D$ as follows
@@
Finally, the visual representation of the RM is the RP of interest. Typically, RPs represent each zero entry ($R_{ij}=0$) of the RM as a white pixel, while each $1$ entry ($R_{ij}=1$) is represented by a black pixel. To ensure that each recurrent point is represented by a unique pixel, the desirable resolution of an RP is $T\times T$ pixels, where $T$ is the maximum iteration time of the dynamics being analyzed. This resolution ensures that the recurrence patterns are accurately represented without overlap, making it easier to visually distinguish between different dynamic behaviors. The optimal choice of resolutions and $T$ is further explored in the appendix A.
II.2 Models
The two dynamical systems studied in this paper are well known nonlinear maps. The first, introduced by Lorenz in the 1960s [22] and popularized by May in 1976 [23], the logistic map, is a standard example of how complex and chaotic behavior can arise from a simple one-dimensional nonlinear equation, with applications ranging from population dynamics, pseudo-random number generators to bifurcation theory [1].
In addition to the logistic map, the Chirikov standard map, an example of a two-dimensional near-integrable Hamiltonian system, was also selected. First conceptualized in the 1970s in the context of plasma physics [24, 25], the standard map exhibits a rich variety of dynamical behaviors, in particular Hamiltonian chaos, making it an ideal subject for this investigation.
@@ II.2.1 The Logistic map
As briefly introduced earlier, the logistic map brings its non-linearity from a simple quadratic term as follows
@@
The logistic map can have several behaviors according to the values taken by $r$, which are briefly described below.
@@
  • (A) $3.569953.0<r<$ 3.56995: Permanent oscillatory behavior is observed, with the number of modes of oscillation depending on $r$;
@@
Figure 1 shows the classical bifurcation diagram along with three distinct RPs, each considering different values of $r$ that properly display the diversity of dynamical behaviors. In agreement with the discussions presented at Sec. II.1 and Appendix A, all RPs from the dynamics of the logistic map were constructed considering a fixed $RR=20\%$, and $T=100$ iterations.
Figure 2: Twenty different dynamical behaviors displayed by their respective RP constructed from the dynamics of the Logistic map. All RPs were build from $T=100$ iterations and fixed $RR=20\%$. All correspondent approximated value of the parameter $r$ is displayed inside each RP.
@@
Since it is important to produce multiple distinct RPs, knowing that neural networks benefit greatly from large volumes of samples, a thorough numerical experiment was conducted to extract as many behaviors from the map as possible. The range $r\in[3.0,4.0]$ was uniformly divided into $10^{4}$ values, generating the same $10^{4}$ different RPs, each differing in the value of $r$ up to the fifth decimal place. The parameters of the dynamics, namely the IC $x_{0}$, and the maximum iteration time $T$ were set to $0.1$ and $100$ respectively. The RP parameters, including the $RR$ and resolution, were set to $20\%$ and $100\times 100$ respectively. A sub-sample of this analysis, consisting of $20$ RPs, is shown in Fig. 2, and all $10^{4}$ images were compressed into a Graphic Interchange Format (GIF) made available in the supplemental material.
It is worth remarking that the initial periodic behaviors, displayed by the first row of RPs in Fig. 2, may exhibit patterns that do not necessarily match the typical appearance of fully periodic RPs [14, 13]. This is because, instead of using a predefined threshold distance $\varepsilon$, these RPs were constructed with a fixed $RR$. Hence, as periodic dynamics presents multiple recurrences, especially for $r\approx 3.0$, the RP is proportionally filled by points until reaching the desired $RR=20\%$. It is also interesting to observe the diversity of patterns within the chaotic regime, followed by an unambiguous periodic behavior produced by the stability islands for $r\approx 3.83$, before returning to chaos as $r$ increases.
@@ II.2.2 The Standard map
In addition to the mixed phase spaces, the lower panels of Fig. 3 also display three RPs associated with orbits A, B and C. These particular trajectories, represented by the black points on their respective phase spaces, are chaotic orbits initiated from a fixed IC $(q_{0},p_{0})=(10^{-9},\pi)$, meaning that they start extremely close to the unstable fixed point. Indeed, at the beginning of their dynamical evolution, the influence of the fixed point is still dominant, as evidenced by the initial black squares in all RPs. Nevertheless, since it is a neighborhood of an unstable fixed point, the chaotic behavior eventually emerges as the orbits lose stability and explore the available chaotic areas.
Figure 4: Twenty different chaotic behaviors displayed by their respective RP constructed from the dynamics of the Standard map. All RPs were build from $T=100$ iterations and fixed $RR=10\%$. All correspondent approximated value of the parameter $k$ is displayed inside each RP.
Figure 5: General architecture of the CNN implemented. Layers indicated in solid colors were used for both dynamics, with the hatched blocks employed solely for the dynamics of the Standard map, and dropout used only for the Logistic map.
@@
It is important to point out that these sets of RP images, exemplified by Figs. 2 and 4, provide effective visual representations of various finite-time dynamical behaviors. In this sense, understanding all the recurrence patterns associated with these RPs is essential to their characterization. To achieve this, a convolutional neural network is trained to effectively obtain the characteristic control parameters associated with these dynamical behaviors.
II.3 CNN and metrics
The architecture of all convolutional neural networks employed in this investigation was defined as shown in Fig. 5, using the structure available at TensorFlow [29]. As for the choice of hyperparameters, the model consists of: 100 batch size; Conv2D(3,3) for each convolutional layer; AveragePooling2D(2) as the pooling layers; ReLu() as the activation function with Adam as chosen optimizer. Root Mean Squared Error (RMSE) and Categorical Cross Entropy are used as loss functions for regression and classification tasks respectively. The training employed 300 maximum epochs, allowing for early stopping with the criteria of interruption being 0.001 and 20 epochs.
In terms of the amount of samples generated that are being used as inputs, the main models had access to $4\times 10^{4}$ unique RPs for the training step, which were composed of $10^{4}$ uniformly chosen values of $r$ and $k$ for four sets of ICs: $x_{0}=0.10,0.25,0.50$  and $0.75$ for the Logistic map; $(q_{0},p_{0})=(\delta\times 10^{-\gamma},\pi)$ where $\gamma=8$ is fixed and $\delta=2,4,6~{}\text{and}~{}8$ for the Standard map. Different values of $\gamma$ were also considered and are further explained in Appendix A.

### v2

# Parameter Inference in Non-linear Dynamical Systems via Recurrence Plots and Convolutional Neural Networks

@@ Abstract
Inferring control parameters in non-linear dynamical systems is an important task in analysing general dynamical behaviours, particularly in the presence of inherently deterministic chaos. Traditional approaches often rely on system-specific models and involve heavily parametrised formulations, which can limit their general applicability. In this study, we present a methodology that employs recurrence plots as structured representations of non-linear trajectories, which are then used to train convolutional neural networks to infer the values of the control parameter associated with the analysed trajectories. We focus on two representative non-linear systems, namely the logistic map and the standard map, and show that our approach enables accurate estimation of the parameters governing their dynamics. When compared to regression models trained directly on raw time-series data, the use of recurrence plots yields significantly more robust results. Although the methodology does not aim to predict future states explicitly, we argue that accurate parameter inference, when combined with predetermined initial conditions, enables the reconstruction of a system’s evolution due to its deterministic nature. These findings highlight the potential of recurrence-based learning frameworks for the automated identification and characterisation of non-linear dynamical behaviours.
@@ I Introduction
Identifying and characterising the diverse dynamical behaviours of non-linear dynamical systems is a challenge, particularly due to the inherent complexity arising from chaotic dynamics [1, 2]. Even in the deterministic scenario, these systems can exhibit multiple distinct dynamics, including periodic, quasi-periodic and chaotic behaviours [3]. Periodic behaviours repeat themselves regularly, whereas quasi-periodic dynamics display structured but non-repeating patterns [4, 5]. Chaotic dynamics, on the other hand, evolve unpredictably over time due to their extreme sensitivity to initial conditions (ICs), despite being governed by fully deterministic rules [6, 7]. Understanding these different behaviours and accurately inferring their governing parameters is essential for gaining deeper insight into the fundamental properties of such systems.
@@
One way to approach such problem is to benefit from the significant developments to neural networks architectures that have been made in the last decade [8], with the introduction of several robust and versatile architectures that allows one to solve high-dimensional partial differential equations (PDEs) [9], identify jets in high-energy physics [10] and more broadly for the accurate identification of images in classification problems [11]. Recently, the relevance of neural networks has been highlighted by the 2024 Nobel Prize in Physics, awarded to Geoffrey E. Hinton and John J. Hopfield. Convolutional Neural Networks (CNN) [12, 13] are a subset of feed-forward neural networks that are widely used in image, text and audio recognition tasks, providing a strong foundation for natural language processing techniques that are now state-of-the-art in artificial intelligence, while remaining relevant due to their effectiveness in many areas. In non-linear dynamical systems, CNNs have also been shown to identify chaotic dynamics when combined with recurrence analysis, achieving over 92% accuracy [14, 15].
Parallel to these advancements, Recurrence Plots (RPs) have become an established approach for visualising and analysing the recurrence of states within dynamical systems [16, 17, 18]. Different dynamical regimes generate distinct visual recurrence patterns, allowing for the qualitative identification of behaviours such as periodic, chaotic, or even stochastic dynamics [19, 20]. Because RPs are effectively binary images, with pixels representing recurrent or non-recurrent states, they are particularly suitable as inputs for CNNs. Although RPs have been widely employed in the detection and analysis of chaotic dynamics [21, 22, 23, 24], their potential for inferring control parameters solely and directly from the recurrence patterns remains relatively unexplored.
Several approaches have been applied to parameter identification and classification tasks in dynamical systems. Traditional techniques such as Kalman filtering [25, 26] have been used in linear and weakly non-linear contexts, providing robust state estimation and parameter tracking capabilities. Similarly, Non-linear Auto-regressive Moving Average with eXogenous inputs (NARMAX) models [27, 28] have proven effective for capturing complex non-linear behaviours, enabling simultaneous model identification and parameter estimation from observed data. A comprehensive overview of more recent developments is provided by Quaranta et al. [29], reviewing state-of-the-art methods for parameter inference in complex systems. Among these, deep learning techniques that include Long Short-Term Memory (LSTM) networks and reservoir computing [30, 31] have emerged as powerful tools for both forecasting and parameter inference in non-linear dynamical systems. In addition, Estebsari et al. [32] implement a NN framework with inputs similar to those we propose, while Kumar and Kostina [33] introduce a Huber loss-based parameter estimation scheme that significantly outperforms classical methods for the Lorenz system. Within this context, our approach differentiates itself by employing RPs as alternative feature spaces for characterising dynamical regimes. Rather than relying solely on raw temporal sequences, we explore how image-based representations of dynamical trajectories enable efficient, data-driven identification of control parameters using CNN-based architectures, also facilitating the analysis of higher-dimensional systems through the RP representation.
Our study introduces a novel approach that combines RPs with CNNs to infer control parameters governing non-linear dynamical behaviours. We apply it to two well-known dynamic systems, the logistic map and the standard map, chosen for their representative non-linear characteristics. By training CNNs on fitting RPs generated from these maps, we assess how accurately the underlying control parameters can be estimated from visual recurrence patterns alone. Additionally, we evaluate the ability of our approach to effectively classify different dynamical regimes. Due to the deterministic nature of these systems, we argue that accurate inference of control parameters, combined with predetermined initial conditions, provides sufficient information to reconstruct their dynamical evolution, even though the method itself does not explicitly predict future states of these systems.
The remainder of this paper is organized as follows. Section II provides detailed descriptions of the methodology, including the construction of RPs, the dynamical characteristics of the logistic and standard maps, and the CNN architecture. Section III presents and discusses our results, analysing the performance of CNNs in the control parameter inference and regime classification tasks. Appendices A and B include additional discussions concerning comparisons with raw time-series regression and considerations for optimal parameter selection. Finally, Section IV summarizes the key findings of this work and suggests possible directions for future research, including potential extensions to more complex and realistic dynamical systems.
@@ II.1 Recurrence Plots
The essence of the proposed methodology relies on the generation of multiple distinct Recurrence Plots (RPs), which serve as fundamental visual representations of the various non-linear dynamical behaviours of interest. A recurrence is typically defined as the return of a trajectory $\bm{x}(t_{j})$ to a neighbourhood of a previously visited state $\bm{x}(t_{i})$, where $t_{i}<t_{j}$. In other words, a state at time $t_{j}$ is considered recurrent if it approximately revisits a past state. Given that this analysis is performed over a finite-time window, we emphasize that $0<t_{i}<t_{j}\leq T$, where $T$ denotes the maximum iteration time of the trajectory.
In the case of one-dimensional dynamical systems, such as the logistic map, constructing an RP requires transforming the scalar time series into a multi-dimensional representation. This is accomplished via time-delay embedding, a technique inspired by Takens’ embedding theorem [34, 35]. The process involves reconstructing the system’s state space by introducing delayed copies of the original time series. Given a scalar sequence $u(t)$, the embedded representation is defined as
$\mathbf{U}(t)=\left[u(t),u(t+\tau),u(t+2\tau),\dots,u(t+(m-1)\tau)\right],$ (1)
where $\tau$ is the time delay, and $m$ is the embedding dimension. For $\tau=1$ and $m=2$, each state in the embedded space consists of the value $u(t)$ and its immediate successor $u(t+1)$, capturing direct temporal relationships between consecutive states, which is sufficient for this methodology. Once the time series is embedded in this higher-dimensional space, an RP is constructed by identifying recurrent states based on a predefined threshold distance, forming the Recurrence Matrix (RM).
For two-dimensional systems, such as the standard map, this additional embedding step is unnecessary. Since the system’s phase space is inherently well-defined—consisting of the generalized pair position-momentum, the RM can be directly extracted by computing distances between trajectory points.
Depending on the model, its control parameters, and the chosen initial conditions, distinct dynamical behaviours may emerge. To ensure a fair and meaningful comparison across chaotic, periodic, and quasi-periodic dynamics, all RPs were constructed using a fixed recurrence rate ($RR$), the percentage of recurrent points in the plot [17]. This normalization allows for an unbiased analysis, ensuring that differences in recurrence structures arise from the underlying dynamics rather than arbitrary choices of threshold distances.
@@
Once the best suited $\varepsilon$ is determined, the aforementioned binary Recurrence Matrix, composed by the elements $R_{i,j}$, $R\in\{0,1\}^{T\times T}$, can be computed in terms of the distance matrix $D$ as follows
@@
Finally, the visual representation of the RM is the RP of interest. Typically, RPs represent each zero entry ($R_{ij}=0$) of the RM as a white pixel, while each $1$ entry ($R_{ij}=1$) is represented by a black pixel. To ensure that each recurrent point is represented by a unique pixel, the desirable resolution of an RP is $T\times T$ pixels, where $T$ is the maximum iteration time of the dynamics being analysed. This resolution ensures that the recurrence patterns are accurately represented without overlap, making it easier to visually distinguish between different dynamical behaviours. The optimal choice of resolutions and $T$ is further explored in the Appendix B.
II.2 Dynamical systems
The two dynamical systems studied in this paper are well known non-linear maps. The first, introduced by Lorenz in the 1960s [36] and popularized by May in 1976 [37], the logistic map, is a classic example of how complex and chaotic behaviour can arise from a simple one-dimensional non-linear equation, with applications ranging from population dynamics, pseudo-random number generators to bifurcation theory [1].
In addition to the logistic map, the Chirikov standard map, an example of a two-dimensional near-integrable Hamiltonian system, was also selected. First conceptualized in the 1970s in the context of plasma physics [38, 39], the map exhibits a rich variety of dynamical behaviours, in particular Hamiltonian chaos, making it an ideal subject for this investigation.
@@ II.2.1 The logistic map
As briefly introduced earlier, the map brings its non-linearity from a simple quadratic term as follows
@@
The logistic map can have several behaviours according to the values taken by $r$, which are briefly described below.
@@
  • (A) $3.569953.0<r<$ 3.56995: Permanent oscillatory behaviour is observed, with the number of modes of oscillation depending on $r$;
@@
Figure 1 shows the classical bifurcation diagram along with three distinct RPs, each considering different values of $r$ that properly display the diversity of dynamical behaviours. In agreement with the discussions presented at Sec. II.1 and Appendix B, all RPs from the dynamics of the logistic map were constructed considering a fixed $RR=20\%$, and $T=100$ iterations.
Figure 2: Twenty different dynamical behaviours displayed by their respective RP constructed from the dynamics of the logistic map. All RPs were build from $T=100$ iterations and fixed $RR=20\%$. All correspondent approximated value of the parameter $r$ is displayed inside each RP.
@@
Since it is important to produce multiple distinct RPs, knowing that neural networks benefit greatly from large volumes of samples, a thorough numerical experiment was conducted to extract as many behaviours from the map as possible. The range $r\in[3.0,4.0]$ was uniformly divided into $10^{4}$ values, generating the same $10^{4}$ different RPs, each differing in the value of $r$ up to the fifth decimal place. The parameters of the dynamics, namely the IC $x_{0}$, and the maximum iteration time $T$ were set to $0.1$ and $100$ respectively. The RP parameters, including the $RR$ and resolution, were set to $20\%$ and $100\times 100$ respectively. A sub-sample of this analysis, consisting of $20$ RPs, is shown in Fig. 2, and all $10^{4}$ images were compressed into a Graphic Interchange Format (GIF) made available in the supplemental material.
It is worth remarking that the initial periodic behaviours, displayed by the first row of RPs in Fig. 2, may exhibit patterns that do not necessarily match the typical appearance of fully periodic RPs [17, 16]. This is because, instead of using a predefined threshold distance $\varepsilon$, these RPs were constructed with a fixed $RR$. Hence, as periodic dynamics presents multiple recurrences, especially for $r\approx 3.0$, the RP is proportionally filled by points until reaching the desired $RR=20\%$. It is also interesting to observe the diversity of patterns within the chaotic regime, followed by an unambiguous periodic behaviour produced by the stability islands for $r\approx 3.83$, before returning to chaos as $r$ increases.
@@ II.2.2 The standard map
In addition to the mixed phase spaces, the lower panels of Fig. 3 also display three RPs associated with orbits A, B and C. These particular trajectories, represented by the black points on their respective phase spaces, are chaotic orbits initiated from a fixed IC $(q_{0},p_{0})=(10^{-9},\pi)$, meaning that they start extremely close to the unstable fixed point. Indeed, at the beginning of their dynamical evolution, the influence of the fixed point is still dominant, as evidenced by the initial black squares in all RPs. Nevertheless, since it is a neighbourhood of an unstable fixed point, the chaotic behaviour eventually emerges as the orbits lose stability and explore the available chaotic areas.
Figure 4: Twenty different chaotic behaviours displayed by their respective RP constructed from the dynamics of the standard map. All RPs were build from $T=100$ iterations and fixed $RR=10\%$. All correspondent approximated value of the parameter $k$ is displayed inside each RP.
Figure 5: General architecture of the CNN implemented. Layers indicated in solid colours were used for both dynamics, with the hatched blocks employed solely for the dynamics of the standard map, and dropout used only for the logistic map.
@@
It is important to point out that these sets of RP images, exemplified by Figs. 2 and 4, provide effective visual representations of various finite-time dynamical behaviours. In this sense, understanding all the recurrence patterns associated with these RPs is essential to their characterization. To achieve this, a convolutional neural network is trained to effectively obtain the characteristic control parameters associated with these dynamical behaviours.
II.3 Convolutional Neural Networks and metrics
The architecture of all convolutional neural networks employed in this investigation was defined as shown in Fig. 5, using the structure available at TensorFlow [43]. As for the choice of hyper-parameters, the model consists of: 100 batch size; Conv2D(3,4) for each convolutional layer; AveragePooling2D(2) as the pooling layers; ReLu() as the activation function with Adam as chosen optimizer. Root Mean Squared Error (RMSE) and Categorical Cross Entropy are used as loss functions for regression and classification tasks respectively. The training employed 300 maximum epochs, allowing for early stopping with the criteria of interruption being 0.001 and 20 epochs. The training set was also split in a 70% to 30% ratio, with the smaller subset being used for validation. The evaluation of validation loss was also used to arrive at the best combinations of the number of layers and parameters to be employed in the following simulations.
In terms of the amount of samples generated that are being used as inputs, the main models had access to $4\times 10^{4}$ unique RPs for the training step, which were composed of $10^{4}$ uniformly chosen values of $r$ and $k$ for four sets of ICs: $x_{0}=0.10,0.25,0.50$  and $0.75$ for the logistic map; $(q_{0},p_{0})=(\delta\times 10^{-\gamma},\pi)$ where $\gamma=8$ is fixed and $\delta=2,4,6~{}\text{and}~{}8$ for the standard map. Different values of $\gamma$ were also considered and are further explained in Appendix B.